# Zeroth-Order Constrained Optimization from a Control Perspective via Feedback Linearization

### v1

@@ 3.2 Theoretical Guarantees on Constraint Satisfaction
$\|h(x)\|\leq H,~~\|J_{h}(x)\|\leq\overline{L}_{h},~~\sigma_{\min}(J_{h}(x))\geq\underline{L}_{h},~~\|D^{2}h(x)\|\leq M,~~\|D^{3}h(x)\|_{\textup{diag}}\leq R.~~\forall~x\in\mathcal{D},$
@@
    $\textstyle D^{3}f(x):\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{m},\quad D^{3}f(x)[u,v,w]=\left.\frac{\partial^{3}}{\partial s\partial t\partial r}f(x+su+tv+rw)\right|_{s=t=r=0}.$
@@
$\textstyle\|D^{2}f(x)\|_{\mathrm{diag}}$ $\textstyle=\sup_{\|u\|=1}\|D^{2}f(x)[u,u]\|,\quad~~\|D^{3}f(x)\|_{\mathrm{diag}}$ $\textstyle=\sup_{\|u\|=1}\|D^{3}f(x)[u,u,u]\|.$
@@
Theorem 1.
Suppose Assumptions 13 hold and $K\succ 0$. Run Algorithm 1 with fresh finite-difference directions at each iteration. Fix $\delta\in(0,1)$ and horizon $T_{G}\in\mathbb{N}$. If the batch size $T_{B}$ and probe radii $r_{1},r_{2}$ satisfy (cf. Appendix Lemma 5)
@@
$\|h(x_{t})]\|$ $\leq\prod_{s=0}^{t-1}(1\!-\!\eta_{s}\lambda_{\min}(K))\|h(x_{0})\|+C_{2}r_{2}^{2}\!\sum_{s=0}^{t-1}\!\prod_{\tau=s\!+\!1}^{t-1}\!(1\!-\!\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}+C_{1}\!\sum_{s=0}^{t-1}\!\prod_{\tau=s\!+\!1}^{t-1}\!(1\!-\!\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}^{2},$
@@
$\|h(x_{t})\|\;\leq\;(1-\eta\lambda_{\min}(K))^{t}\,\|h(x_{0})\|~+~\frac{C_{2}r_{2}^{2}}{\lambda_{\min}(K)}~+~\frac{C_{1}\eta}{\lambda_{\min}(K)}.$ (12)
@@
$\|h(x_{t})\|\;\leq\;e^{-\eta\lambda_{\min}(K)(\sqrt{t}-1)}\,\|h(x_{0})\|~+~\frac{2e\,C_{2}r_{2}^{2}}{\lambda_{\min}(K)}~+~\frac{C_{1}\eta\,e^{2-\eta\sqrt{t}}}{\lambda_{\min}(K)}~+~\frac{2e\,C_{1}\eta}{\lambda_{\min}(K)\sqrt{t+1}}.$ (13)
Remark 1 (Interpretation of Constraint Violation Bound).
We now unpack the meaning of the bound in Eq. 12. The first term, $(1-\eta{\lambda_{\min}}(K))^{t}\|h(x_{0})\|$, decays exponentially in $t$ to zero. This reflects the core effect of the FL design: in the absence of estimation or discretization errors, the constraint dynamics reduce to a simple stable linear system, driving violations to zero at a geometric rate. In this sense, ZOFL inherits the strong feasibility guarantees of first-order FL.
@@
The third term, $\frac{C_{1}\eta}{{\lambda_{\min}}(K)}\sim O(\eta)$, comes from higher-order terms in the Taylor expansion of the constraint dynamics. Unlike the approximation error, this residual is intrinsic to the Euler discretization used in ZOFL: even with perfect gradient and Jacobian information, a fixed step size $\eta$ produces a non-vanishing bias. This is the main bottleneck for achieving exact feasibility under constant step sizes.
To mitigate this discretization bias, one can use a diminishing schedule as in Eq. 13 In this case, the residual terms vanish asymptotically, and constraint violations eventually disappear. The trade-off is that the ideal contraction term slows down: instead of exponential decay, the dominant term becomes $e^{-\eta\lambda_{\min}(K)(\sqrt{t}-1)}\|h(x_{0})\|$, which decreases subexponentially in $t$. This mirrors a common theme in stochastic optimization: stronger asymptotic guarantees are possible, but only at the cost of slower transient progress.
@@ 4 Extension to Inequality-constrained Setting
$\min_{x\in\mathbb{R}^{n}}f(x)\quad\text{s.t. }h(x)\leq 0.$ (14)
$\hskip-68.0pt\textup{The KKT conditions are}~~-\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)\leq 0,\quad\lambda\geq 0,\quad\lambda^{\top}h(x)=0.$ (15)
@@
$\begin{split}\textstyle x_{t+1}-x_{t}=-\eta_{t}\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right),\quad y_{t}=h(x_{t}),\quad\lambda_{t}\geq 0.\end{split}$ (16)
Compared with the equality case, the difficulty lies in enforcing the non-negativity of multipliers and the complementary slackness condition $\lambda^{\top}h(x)=0$. In (Zhang et al., 2025), this is achieved by designing a more intricate FL controller:
FO-FL for Inequality-Constrained Optimization $\begin{split}x_{t+1}-x_{t}&=-\eta_{t}\!\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right),\\ \lambda_{t}&=\operatorname*{arg\,min}_{\lambda\geq 0}\Big\{\tfrac{1}{2}\lambda^{\top}J_{h}(x_{t})J_{h}(x_{t})^{\top}\lambda+\lambda^{\top}\!\big(J_{h}(x_{t})\nabla f(x_{t})-Kh(x_{t})\big)\Big\}.\end{split}$ (17)
@@
$\textstyle\lambda_{t}=\operatorname*{arg\,min}_{\lambda\geq 0}\Big\{\tfrac{1}{2}\lambda^{\top}J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\lambda+\lambda^{\top}\!\big(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})\big)\Big\}.$ (18)
@@
$J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}+J_{h}(x_{t})\widetilde{\nabla}f(x_{t})=Kh(x_{t})+s,\quad s^{\top}\lambda_{t}=0,\quad s\geq 0,\quad\lambda_{t}\geq 0.$ (19)
@@
    $\begin{split}{G_{h}}\lambda+{G_{f}}=Kh(x_{t})+s,~~~s^{\top}\lambda=0,~~~s\geq 0,~~~\lambda\geq 0\end{split}$
@@
Theoretical guarantees.
We now state the main feasibility result. The proof follows the same high-level structure as Theorem 1 but requires sharper bounds on the error terms due to the nonsmooth projection step.
Theorem 2 (Feasibility with Inequality Constraints).
Under Assumption 1, 2 and 3, suppose $T_{B}$ and $r_{1},r_{2}$ are chosen as in Theorem 1. Then with probability at least $1-\delta$, the ZOFL algorithm for inequality constraints (Algorithm 2) satisfies
$\|[h(x_{t})]_{+}\|$ $\leq\prod_{s=0}^{t-1}(1\!-\!\eta_{s}\lambda_{\min}(K))\|[h(x_{0})]_{+}\|+C_{2}r_{2}^{2}\!\sum_{s=0}^{t-1}\!\prod_{\tau=s\!+\!1}^{t-1}\!(1\!-\!\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}+C_{1}\!\sum_{s=0}^{t-1}\!\prod_{\tau=s\!+\!1}^{t-1}\!(1\!-\!\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}^{2},$
@@
$C_{1}=M\!\left(nL_{f}+\tfrac{64n\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right),\qquad C_{2}=n^{2}\overline{L}_{h}^{2}R\!\left(\tfrac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\tfrac{64L_{f}}{\underline{L}_{h}^{2}}\right).$
@@
For diminishing step $\eta_{t}=\eta/\sqrt{t}$, the bound improves asymptotically as in Theorem 1.
The structure of the bound mirrors the equality-constrained case: exponential contraction toward feasibility, plus two residual terms accounting for zeroth-order approximation and discretization. The detailed interpretation in 1 applies here as well, with the caveat that violations are measured via $h(x_{t})]_{+}$ rather than $h(x_{t})$.
@@ 5 Exploring midpoint methods for zeroth-order optimization
In Remark 1, we pointed out that discretization error is a major bottleneck in controlling constraint violation. This error arises from approximating $h(x_{t+1})-h(x_{t})$ using only the first-order term of the Taylor expansion, leading to an $O(\eta)$ residual. A natural question, then, is whether more accurate numerical schemes can reduce this error. Motivated by this, we introduce the midpoint method from numerical analysis (cf. Süli and Mayers (2003)), which achieves a discretization error of $O(\eta^{2})$, and develop the midpoint variant of ZOFL (Algorithm 3). Our experiments (Figures 3(a) and 3(b)) demonstrate that this variant achieves improved constraint satisfaction compared to standard ZOFL. However, ZOFL-midpoint requires twice as many function evaluations per iteration, highlighting a trade-off between accuracy and sample efficiency. We further conjecture that the constraint violation bound under the midpoint method scales as $O(\eta^{2})$, and leave a rigorous proof of this property as an open question.
@@ 6 Numerical Validations
We implement the ZOFL and ZOFL-midpoint algorithms (Algorithm 1 and 3) and compare it with the ZO-baseline method (equation 8) along with other baseline algorithms in zeroth-order constrained optimization, namely SZO-ConEx (Nguyen and Balasubramanian (2022)) and ZOGDA Liu et al. (2020).
@@
Inequality Constrained. We tested our algorithm on learning an efficient controller for building thermal regulation. We assume that the thermal dynamics to be a linear RC model (Zhang et al. (2016); Li et al. (2022a)) $x_{t+1}=Ax_{t}+Bu_{t}+d,$ where $x_{t}=\{x_{1,t},x_{2,t},\dots,x_{n,t}\}\in\mathbb{R}^{n}$ represents the temperature in each building at time step $t$, $u_{t}=\{u_{1,t},u_{2,t},\dots,u_{n,t}\}$ is the thermal power injection and $d$ is the disturbances. We consider the controller  $u_{i,t}=k_{i}x_{i,t}+b_{i}$  and the optimization problem is given by optimizing the control parameters: $K=\{k_{i}\}_{i=1}^{n},b=\{b_{i}\}_{i=1}^{n}$ to minimize the thermal energy subject to the thermal comfort constraint:

### v2

@@ 3.2 Theoretical Guarantees on Constraint Satisfaction
$\|h(x)\|\!\leq\!H,~~\|J_{h}(x)\|\!\leq\!\overline{L}_{h},~~\sigma_{\!\min}(J_{h}(x))\!\geq\!\underline{L}_{h},~~\|D^{2}h(x)\|\!\leq\!M,~~\|D^{3}h(x)\|_{\textup{diag}}\!\leq\!R$
@@
    $\textstyle D^{3}f(x):\mathbb{R}^{n}\!\times\!\mathbb{R}^{n}\!\times\!\mathbb{R}^{n}\!\to\!\mathbb{R}^{m},~D^{3}f(x)[u,v,w]\!=\!\left.\frac{\partial^{3}}{\partial s\partial t\partial r}f(x\!+\!su\!+\!tv\!+\!rw)\right|_{s=t=r=0}\!\!.$
@@
$\textstyle\|D^{2}f(x)\|_{\mathrm{diag}}\!=\!\sup_{\|u\|\!=\!1}\!\|D^{2}f(x)[u,u]\|,~\|D^{3}f(x)\|_{\mathrm{diag}}\!=\!\sup_{\|u\|=1}\|D^{3}f(x)[u,u,u]\|.$
@@
Theorem 3.1.
Suppose Assumptions 13 hold and $K\succ 0$. Run Algorithm 1 with $u_{i}$’s n (7) drawn i.i.d. from the unit sphere. Fix $\delta\in(0,1)$ and horizon $T_{G}\in\mathbb{N}$. If the batch size $T_{B}$ and probe radii $r_{1},r_{2}$ satisfy (cf. Appendix Lemma E.1)
@@
$\textstyle\|h(x_{t})]\|\leq\prod_{s=0}^{t-1}(1-\eta_{s}\lambda_{\min}(K))\|h(x_{0})\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}$
$\textstyle+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}^{2},$
@@
(12) $\|h(x_{t})\|\;\leq\;(1-\eta\lambda_{\min}(K))^{t}\,\|h(x_{0})\|~+~\frac{C_{2}r_{2}^{2}}{\lambda_{\min}(K)}~+~\frac{C_{1}\eta}{\lambda_{\min}(K)}.$
@@
(13) $\|h(x_{t})\|\;\leq\;e^{-\eta\lambda_{\min}(K)(\sqrt{t}-1)}\,\|h(x_{0})\|~+~\frac{2e\,C_{2}r_{2}^{2}}{\lambda_{\min}(K)}~+~\frac{C_{1}\eta\,e^{2-\eta\sqrt{t}}}{\lambda_{\min}(K)}~+~\frac{2e\,C_{1}\eta}{\lambda_{\min}(K)\sqrt{t+1}}.$
Remark 1. (Interpretation of Constraint Violation Bound) We now unpack the meaning of the bound in Eq. 12. The first term, $(1-\eta{\lambda_{\min}}(K))^{t}\|h(x_{0})\|$, decays exponentially in $t$ to zero. This reflects the core effect of the FL design: in the absence of estimation or discretization errors, the constraint dynamics reduce to a simple stable linear system, driving violations to zero at a geometric rate. In this sense, ZOFL inherits the strong feasibility guarantees of first-order FL.
@@
The third term, $\frac{C_{1}\eta}{{\lambda_{\min}}(K)}\sim O(\eta)$, comes from higher-order terms in the Taylor expansion of the constraint dynamics. Unlike the approximation error, this residual is intrinsic to the Euler discretization used in ZOFL, where we approximate $h(x_{t+1})-h(x_{t})$ with the first order Taylor expansion $J_{h}(x_{t})(x_{t+1}-x_{t})$ (see (9)). Thus a fixed step size $\eta$ produces a non-vanishing bias. This is the main bottleneck for achieving exact feasibility under constant step sizes.
To mitigate this discretization bias, one can use a diminishing schedule as in Eq. 13. In this case, the residual terms vanish asymptotically, and constraint violations eventually disappear. The trade-off is that the ideal contraction term slows down: instead of exponential decay, the dominant term becomes $e^{-\eta\lambda_{\min}(K)(\sqrt{t}-1)}\|h(x_{0})\|$, which decreases subexponentially in $t$. This mirrors a common theme in stochastic optimization: stronger asymptotic guarantees are possible, but at the cost of slower transient progress.
@@
3.3 Global Convergence
While Theorem 3.1 establishes guarantees for constraint satisfaction, it does not address the convergence to a KKT point. In this section, we will establish global convergence guarantees for the equality constrained setting (1). We will make the following additional assumption
Assumption 4.
The objective function is $M_{f}$-smooth, i.e., $\|\nabla^{2}f(x)\|\leq M_{f}$ and its third order derivative satisfies $\|D^{3}f(x)\|_{\textup{diag}}\leq R$.
Further in order for the analysis to carry through, we add an additional line of code between Step 2.1 and 2.2 in Algorithm 1 to check if $\sigma_{\min}(G_{h})\geq\sigma>0$, where $\sigma$ is some constant that we choose such that $\sigma>\frac{\underline{L}_{h}}{64}$, and if this condition is not met, we reject the sampled $u_{i}$’s and redo the sampling until the condition is met.
Theorem 3.2.
Under Assumption 1, 2, 3, and 4, by running the above described modified Algorithm 1, for $\eta_{t}\!=\!\eta\!\leq\!\frac{1}{6n(M_{f}+\tau M)}$ where $\tau\!=\!64(T_{B}L_{f}+H)\overline{L}_{h}\underline{L}_{h}^{-2}\frac{\|K\|}{{\lambda_{\min}}(K)}$, we have that
$\textstyle\liminf_{t\to+\infty}\left(\|\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda^{\star}(x_{t})\|^{2}+\|h(x_{t})\|_{1}\right)\leq{\frac{64T_{B}\epsilon}{\eta n}},$
where $\lambda^{\star}(x_{t}):=(J_{h}(x_{t})J_{h}(x_{t})^{\top})^{-1}J_{h}(x_{t})\nabla f(x_{t}),\epsilon:=n\overline{L}_{h}\underline{L}_{h}^{-2}Rr_{1}^{2}+C_{2}r_{2}^{2}$ ($C_{2}$ defined as in Theorem 3.1).
The proof of the theorem is in Appendix C. We would like to remark that the stepsize $\eta$ in the theorem is generally too small and the algorithm converges too slowly. Thus in practice, it is better to adopt linesearch methods to determine the stepsize adaptively. We would like to leave it as future work about more efficient linesearch design.
@@ 4 Extension to Inequality-constrained Setting
(14) $\min_{x\in\mathbb{R}^{n}}f(x)\quad\text{s.t. }h(x)\leq 0.$
The KKT conditions are
(15) $\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)\leq 0,\quad\lambda\geq 0,\quad\lambda^{\top}h(x)=0.$
@@
(16) $\begin{split}\textstyle x_{t+1}-x_{t}=-\eta_{t}\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right),\quad y_{t}=h(x_{t}),\quad\lambda_{t}\geq 0.\end{split}$
Compared with the equality case, the difficulty lies in enforcing the non-negativity of multipliers and the complementary slackness condition $\lambda^{\top}h(x)=0$. In [zhang2025constrainedoptimizationcontrolperspective], this is achieved by designing a more intricate FL controller:
FO-FL for Inequality-Constrained Optimization (17) $\begin{split}&\textstyle x_{t+1}-x_{t}=-\eta_{t}\!\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right),\\ &\textstyle\lambda_{t}=\operatorname*{arg\,min}_{\lambda\geq 0}\Big\{\tfrac{1}{2}\lambda^{\top}J_{h}(x_{t})J_{h}(x_{t})^{\top}\lambda\!+\!\lambda^{\top}\!\big(J_{h}(x_{t})\nabla f(x_{t})\!-\!Kh(x_{t})\big)\!\Big\}.\end{split}$
@@
(18) $\textstyle\lambda_{t}=\operatorname*{arg\,min}_{\lambda\geq 0}\Big\{\tfrac{1}{2}\lambda^{\top}J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\lambda+\lambda^{\top}\!\big(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})\big)\Big\}.$
@@
(19) $J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}+J_{h}(x_{t})\widetilde{\nabla}f(x_{t})=Kh(x_{t})+s,~~s^{\top}\lambda_{t}=0,~~s\geq 0,~~\lambda_{t}\geq 0.$
@@
    $\begin{split}\textstyle{G_{h}}\lambda+{G_{f}}=Kh(x_{t})+s,~~~s^{\top}\lambda=0,~~~s\geq 0,~~~\lambda\geq 0\end{split}$
@@
Theoretical guarantees.  We now state the main feasibility result. The proof follows the same high-level structure as Theorem 3.1 but requires sharper bounds on the error terms due to the nonsmooth projection step.
Theorem 4.1 (Feasibility with Inequality Constraints).
Under Assumption 1, 2 and 3, suppose $T_{B}$ and $r_{1},r_{2}$ are chosen as in Theorem 3.1. Then with probability at least $1-\delta$, the ZOFL algorithm for inequality constraints (Algorithm 2) satisfies
$\textstyle\|[h(x_{t})]_{+}\|\leq\prod_{s=0}^{t-1}(1\!-\!\eta_{s}\lambda_{\min}(K))\|[h(x_{0})]_{+}\|+C_{2}r_{2}^{2}\!\sum_{s=0}^{t-1}\!\prod_{\tau=s\!+\!1}^{t-1}\!(1\!-\!\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}$
$\textstyle+C_{1}\!\sum_{s=0}^{t-1}\!\prod_{\tau=s\!+\!1}^{t-1}\!(1\!-\!\eta_{\tau}\lambda_{\min}(K))\,\eta_{s}^{2},$
@@
$C_{1}=M\!\left(nL_{f}+\tfrac{64n\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right),~~C_{2}=n^{2}\overline{L}_{h}^{2}R\!\left(\tfrac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\tfrac{64L_{f}}{\underline{L}_{h}^{2}}\right).$
@@
For diminishing step $\eta_{t}=\eta/\sqrt{t}$, the bound improves asymptotically as in Theorem 3.1.
The structure of the bound mirrors the equality-constrained case: exponential contraction toward feasibility, plus two residual terms accounting for zeroth-order approximation and discretization. The detailed interpretation in Remark 3.1 applies here as well, with the caveat that violations are measured via $h(x_{t})]_{+}$ rather than $h(x_{t})$.
@@ 5 Exploring midpoint methods for zeroth-order optimization
In Remark 3.1, we pointed out that discretization error is a major bottleneck in controlling constraint violation. This error arises from approximating $h(x_{t+1})-h(x_{t})$ using only the first-order term of the Taylor expansion, leading to an $O(\eta^{2})$ residual. A natural question, then, is whether more accurate numerical schemes can reduce this error. Motivated by this, we introduce the midpoint method from numerical analysis (cf. [suli2003introduction]), which achieves a discretization error of $O(\eta^{3})$, and develop the midpoint variant of ZOFL (Algorithm 3). Our experiments (Figures 3(a) and 3(b)) demonstrate that this variant achieves improved constraint satisfaction compared to standard ZOFL. However, ZOFL-midpoint requires twice as many function evaluations per iteration, highlighting a trade-off between accuracy and sample efficiency. We further conjecture that the constraint violation bound under the midpoint method scales as $O(\eta^{2})$, and leave a rigorous proof of this property as an open question.
@@ 6 Numerical Validations
We implement the ZOFL and ZOFL-midpoint algorithms (Algorithm 1 and 3) and compare it with the ZO-baseline method ((8)) along with other baseline algorithms in zeroth-order constrained optimization, namely SZO-ConEx ([nguyen_stochastic_2022]) and ZOGDA [liu_min-max_2020].
@@
Inequality Constrained. We tested our algorithm on learning an efficient controller for building thermal regulation. We assume that the thermal dynamics to be a linear RC model ([zhang2016decentralized, Li2022Distributed]) $x_{t+1}=Ax_{t}+Bu_{t}+d,$ where $x_{t}=\{x_{1,t},x_{2,t},\dots,x_{n,t}\}\in\mathbb{R}^{n}$ represents the temperature in each building at time step $t$, $u_{t}=\{u_{1,t},u_{2,t},\dots,u_{n,t}\}$ is the thermal power injection and $d$ is the disturbances. We consider the controller  $u_{i,t}=k_{i}x_{i,t}+b_{i}$  and the optimization problem is given by optimizing the control parameters: $K=\{k_{i}\}_{i=1}^{n},b=\{b_{i}\}_{i=1}^{n}$ to minimize the thermal energy subject to the thermal comfort constraint: